OpenStudy (rose4825):

A right cone has a base diameter that is 6 cm and a slant height that is 27 cm. What is the surface area of the cone? 81π cm2 84π cm2 162π cm2 90π cm2 Please explain.

OpenStudy (rose4825):

@jim_thompson5910 @jhonyy9

OpenStudy (rose4825):

I know how to do it, but I cannot get an answer with pi left out

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

why not get it witout pi - just dont use in your calcules this value of pi of 3,14 and what you get this number will use to answer togheter with pi

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (rose4825):

Oh okay... so it's 81pi^2

jhonyy9 (jhonyy9):

pi squared not is right the unit measure you need writing squared - ok. ?

OpenStudy (rose4825):

Lol the first one, then?

OpenStudy (rose4825):

That's what I meant

jhonyy9 (jhonyy9):

so will be 81pi cm^2

OpenStudy (rose4825):

Okay

jhonyy9 (jhonyy9):

do yu understand it now sure right ?

OpenStudy (rose4825):

Yeah I get it. I knew how to do it, but I was confused on the pi part.

OpenStudy (rose4825):

Can you quickly check this one: http://prntscr.com/bg2f29

jimthompson5910 (jim_thompson5910):

you made a typo somewhere. The answer isn't 81pi

jimthompson5910 (jim_thompson5910):

I'm referring to the first question at the top

jimthompson5910 (jim_thompson5910):

Surface area of a cone \[\Large SA = \pi*r*(r+s)\] In this case r = 3 is your radius s = 27 is the slant height \[\Large SA = \pi*r*(r+s)\] \[\Large SA = \pi*3*(3+27)\] \[\Large SA = \pi*3*(30)\] \[\Large SA = 90\pi\]

OpenStudy (rose4825):

Wait, what?

OpenStudy (rose4825):

The equation for the surface area of the cone with slant height is A=πrl+πr^2

jimthompson5910 (jim_thompson5910):

r = 3 is the radius L = 27 is the slant height \[\Large SA = \pi*r*L+\pi*r^2\] \[\Large SA = \pi*3*27+\pi*3^2\] \[\Large SA = \pi*81+\pi*9\] \[\Large SA = 81\pi+9\pi\] \[\Large SA = 90\pi\]

jimthompson5910 (jim_thompson5910):

The formula I used is in a different form but it's equivalent to your formula

jhonyy9 (jhonyy9):

JIm sorry but i think Rose mean total area and you wrote surface area what mean lateral area - than i m right ???

OpenStudy (rose4825):

Oh... okay. @jhonyy9 I need the surface area of the cone

OpenStudy (rose4825):

@jim_thompson5910 Is 90pi cm^2 correct?

jimthompson5910 (jim_thompson5910):

`Is 90pi cm^2 correct?` yes @rose4825

OpenStudy (rose4825):

@jim_thompson5910 Can you check this one? http://prntscr.com/bg2f29

jimthompson5910 (jim_thompson5910):

ok checking now

jimthompson5910 (jim_thompson5910):

9 is correct